The first term of an AP is 6 and the fifth term is 12. The progression has n terms and the sum of all the terms is 90.
What is the value of n? (how many terms)

Answer by mananth(16946)

The first term of an AP is 6 and the fifth term is 12. The progression has n terms and the sum of all the terms is 90.
What is the value of n? (how many terms)
Tn = a+(n-1)d
a=6 ,
T5 = 12
12=6+(5-1)d
12=6+4d
12-6=4d
6=4d
d=3/2
Tn = 90
90=6+(n-1)*1.5
90-6= 1.5n -1.5
84+1.5=1.5n
85.5/1.5 = n
57 = n